Calculate the cfu/ml using the formula

cfu/ml = # colonies x dilution factor x plating factor (10)

Temp. (OC) Organism #Colonies Dilution Staphylococcus aureus 43 15 1:1000000 51 1:10000 1:100 53 6
Answer #1

1. CFU = 15 × 10^6 × 10 = 15 × 10^7 CFU per ml

2. CFU = 8 × 10^4 × 10 = 8 × 10^5 CFU per ml

3. CFU = 6 × 10^2 = 600 CFU per ml

Dilution factor = 1/dilution

So, if dilution is 1/100, then dilution factor will be 100.

Dilution factor = final volume / initial volume

Please rate.